OpenStudy (anonymous):

find three consecutive integers such that twice the greatest integer is 2 less than 3 times the least integer

OpenStudy (bigbosssaint21):

6,7,8

OpenStudy (texaschic101):

3 consecutive integers means 1 right after the other. a = lowest integer a + 1 = middle integer a + 2 = largest integer 2(a + 2) = 3(a) - 2 2a + 4 = 3a - 2 4 + 2 = 3a - 2a 6 = a a + 1 = 7 a + 2 = 8 it is 6,7,8

OpenStudy (anonymous):

thanks for not only helping me but u explained how to do it ...now i understand @texaschic101

OpenStudy (texaschic101):

no problem :)